OpenStudy (anonymous):

hey anyone knowhow to calculate newtons laws on angular motion ? Q, what torque is reqiired to produce an angular acceleration of 40 rads/s sq in a wheel that has a moment of inertia of 400gk-m sq ?

OpenStudy (anonymous):

You have: \[\alpha=30\frac{1}{s^2}\quad I=400\ kg\cdot m^2\] For angular acceleration alpha, torque T and moment of inertia I we have: \[ T=I\cdot \alpha \] and the answer to your question is: \[ T=400kg\cdot m^2\cdot 30\frac{1}{s^2}=12000\ Nm. \] To study more, see: http://en.wikipedia.org/wiki/Uniform_circular_motion http://en.wikipedia.org/wiki/Moment_of_inertia

OpenStudy (anonymous):

a is actually 40, which makes the answer 16000. The rest is right :D